Ask your own question, for FREE!
Mathematics 10 Online
OpenStudy (anonymous):

If you buy a school raffle ticket for $1, your overall probability of winning at least one of the prizes is 0.03. You can improve your odds of winning something by buying two tickets. What is the probability that one or the other of your tickets will win something?(Round your answer to the nearest ten-thousandth.)

OpenStudy (anonymous):

In first case: Chance = 3/100 = 1/(100/3) Second case: Chance = 2/(100/3+1) = 0.58

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!